The Government of Pakistan is emphasizing on the consumption of indigenous and environmentally clean energy generation resources. Alternative Energy Development Board (AEDB) is established by the government, responsible for the growth of renewable energy sector; the advancement of alternative and renewable technologies is among one of the top priorities of the government. Several initiatives have been taken to create favorable environment for the sustainable growth of Alternative Renewable Energy (ARE) Sector in Pakistan in order to control and make potential use of indigenous renewable energy resources. Suppose the quantity demanded and quantity supplied functions for renewable energy resources are as given below: Qd = 20,000 - 3P Qs = 15,000 + 2P Where ‘P’ is the price in rupees of a renewable energy resources, ‘Qd’ is quantity demanded for renewable energy resources, and ‘Qs’ is quantity supplied for renewable energy resources.


1
Expert's answer
2021-12-06T16:59:17-0500

Solution

"Qd = 20,000 - 3P"

"Qs = 15,000 + 2P"


Determination of Market Equilibrium Price and Quantity of Renewable Energy Resources:

At Equilibrium Point "Q_d=Q_s"

So,

"20,000 - 3P=15,000 + 2P"

"-5P=-5000"

"P=1000"

So, equilibrium price is 1000


Substitute P into either "Qd" or "Q_s"

"Q=20,000-3(1000)=20,000-3,000=17,000" units

So, equilibrium quantity is 17,000 units


Therefore, Pakstan should set price for renewable energy resources in Pakistan market at 1000 to ensure fair supply and demand in the market. This will reduce excessive demand and supply in the market thus fostering growth of renewable energy sector.